# Federated Sketching LoRA: On-Device Collaborative Fine-Tuning of Large Language Models

### v2

@@ 3.2 Sparsity in the Gradients
In this subsection, we analyze the gradient structure of LoRA modules and highlight the gradients’ sparsity properties under a given sketching matrix. To begin, we present the gradient expressions for the LoRA modules $\bf B$ and $\bf A$ in the following lemma. The proof is provided in Appendix F.2.
@@
In particular, a random-$k$ diagonal sketching matrix selectively samples $k$ rows or columns of a matrix through left product or right product, respectively. With ${\bf S}$ being a random-$k$ diagonal matrix, the gradients of $\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$ with respect to LoRA modules $\bf B$ and $\bf A$, as shown in (3), naturally become structurally sparse matrices. This sparsity reduces the computational and memory overhead during training, allowing for faster gradient computation and parameter updates. Additionally, sparse training enables better scalability across distributed devices by reducing communication costs, as only the non-zero elements need to be transmitted.
@@
A key advantage of our formulation is its flexible control over the sparsity level of local gradients, achieved by configuring the parameter $k_{i}$ of the sketching matrix set $\mathcal{S}_{i}=\mathcal{S}(r,k_{i})$. This mechanism allows each device to tailor its local updates according to its communication and computation resource constraints, ensuring efficient and scalable fine-tuning in heterogeneous federated systems. Lowering $k_{i}$ helps resource-constrained devices reduce computation and communication overhead, while more capable devices can increase $k_{i}$ to conduct more informative local updates. Additionally, the distinction in sparsity level control between the proposed FSLoRA and the FedBCGD algorithm [26] is elaborated in Appendix B.
@@ 3.3 FSLoRA Algorithm
Based on the formulation in (2), we propose a resource-adaptive algorithm termed FSLoRA for collaborative on-device fine-tuning. FSLoRA allows each device to update submatrices of the original modules $\bf{B}$ and $\bf{A}$ in each round. The server maintains a pair of global LoRA modules $\mathbf{B}$ and $\mathbf{A}$ and periodically updates them by aggregating sparse local updates received from distributed devices. Specifically, the procedure of FSLoRA at each round $t$ is detailed below.
  • The server begins by sampling sketching matrices $\{{\bf S}_{i}^{t}\sim\mathcal{S}_{i}\}_{i=1}^{N}$ for all devices, where $\mathcal{S}_{i}$ represents the set of possible sketching matrices for device $i$. These sketches are then sent to the corresponding devices. Additionally, the server broadcasts the current global LoRA modules $[\mathbf{B}^{t};\mathbf{A}^{t}]$ to all devices. Note that the communication load introduced by transmitting the sketching matrix is negligible compared to global LoRA modules, as it involves only binary sketching indices (i.e., the diagonal elements of the sketching matrix); see Appendix A for details.
  • Devices perform local fine-tuning using sketch ${\bf S}_{i}^{t}$. Specifically, guided by sketching matrix ${\bf S}_{i}^{t}$, the update at device $i$ during the $h$-th iteration of the $t$-th round is given by:
@@
    where $\gamma$ denotes the learning rate and $[{\Delta}{\bf B}_{i}^{t,h};{\Delta}{\bf A}_{i}^{t,h}]$ is a shorthand representation for:
@@
    The update direction in (4) corresponds to the negative stochastic gradient of $\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$ with respect to $[{\bf B};{\bf A}]$ for a given sketch ${\bf S}_{i}^{t}$, as established in Lemma 3.2. The total update for device $i$ during one round of training, consisting of $H$ local steps, can be expressed as follows:
@@
    From the above equation, we see that only the columns of $\mathbf{B}$ and the rows of $\mathbf{A}$ corresponding to the nonzero entries of $\mathbf{S}_{i}^{t}$ are updated during the $t$-th round at device $i$. In essence, $\mathbf{S}_{i}^{t}$ selectively activates specific columns of $\mathbf{B}$ and rows of $\mathbf{A}$ for each round. Afterward, devices transmit these nonzero columns and rows of the sparse model updates to the server.
@@
The above procedure is repeated for $t=0,1,\ldots,T\!-\!1$ across $T$ rounds. Algorithm 1 summarizes the overall process of FSLoRA.
@@
Existing works on federated LoRA primarily adopt two aggregation strategies: (1) aggregating the LoRA modules as $[\mathbf{B};\mathbf{A}]$ (e.g., vanilla Federated LoRA [43]), and (2) aggregating the product $\mathbf{B}\mathbf{A}$ (e.g., FlexLoRA [1]). Both methods have demonstrated effectiveness, as evidenced by their promising performance in prior studies. In this work, we adopt the former, as it introduces minimal overhead and retains the simplicity of LoRA. Additionally, we establish the convergence of Algorithm 1 under this aggregation choice, as shown in Section 4.
@@
The proposed FSLoRA introduces no additional operations compared to the vanilla Federated LoRA [43], resulting in minimal overhead for both the server and the devices relative to other heterogeneous LoRA baselines [1, 37]. A more detailed comparison of computation, memory, and communication is provided in Appendix A.
@@
2:     Server samples sketching matrices $\{{\bf S}_{i}^{t}\sim\mathcal{S}_{i}\}_{i=1}^{N}$ and sends them back to the devices
3:     Server broadcasts the current global LoRA modules to the devices
@@
5:        Devices update the local LoRA modules via (4)
@@
7:     Devices upload the non-zero columns of $({\bf B}_{i}^{t,H}\!-\!{\bf B}_{i}^{t,0})$ and the non-zero rows $({\bf A}_{i}^{t,H}\!-\!{\bf A}_{i}^{t,0})$ to the server
@@ 3.4 Comparison with Communication Compression
Although both the sketching approach in FSLoRA and communication compression [22] reduce communication overhead, the sketching approach fundamentally differs from traditional compression techniques. Compression methods focus solely on reducing the transmission load, leaving the gradient computation and model updates unchanged from the vanilla Federated LoRA. FSLoRA goes beyond communication savings by also reducing gradient computation and model update overhead through sparse training. Notably, these two methods are orthogonal and can be combined to achieve greater efficiency. Specifically, the compression can be applied to the transmission of non-zero columns of $\mathbf{B}$ and the non-zero rows of $\mathbf{A}$ in FSLoRA to further enhance communication efficiency. We demonstrate the effectiveness of this combination in Appendix E.2.
@@ 4 Analysis
In this section, we analyze the convergence of the proposed FSLoRA algorithm. We show that the iterate sequence generated by FSLoRA algorithm converges to a stationary point of the function (2). Our analysis relies on the following notations.
[…]
@@
Assumptions 4.1 and 4.2 are standard in stochastic optimization [8, 12], while Assumption 4.3 is commonly used in distributed optimization [11, 40] to characterize data heterogeneity. Building on these assumptions, we analyze the convergence behavior of FSLoRA. Our main results are summarized in the following theorem.
@@
Suppose that Assumptions 4.1-4.3 hold and the learning rate satisfies $\gamma\leq\min\{\frac{N}{24\rho(c_{h}+1)H\bar{L}},\frac{1}{8\sqrt{\widetilde{L}L(\rho+1)(c_{h}+1)}H}\}$. Then the iterates $\{{\bf X}^{t}\}_{t=0}^{T-1}$ generated by FSLoRA satisfy
where $\sigma_{\rho}^{2}=\sigma^{2}+3(\rho+1)\sigma_{h}^{2}$, $\bar{L}=\left(\frac{1}{N}\sum_{i=1}^{N}\frac{r}{k_{i}}\right)L$, $\widetilde{L}=\left(\frac{1}{N}\sum_{i=1}^{N}\frac{r^{2}}{k^{2}_{i}}\right)L$, and $f^{*}$ denotes the lower bound of $f^{\mathcal{S}}({\bf X})$.
Technical highlights of Theorem 4.4: A key step in the proof of Theorem 4.4 is characterizing the impact of sketching mechanism on the optimization landscape. Our analysis reveals how the sketching operation modifies the smoothness properties of the objective, introducing scaled smoothness constants, $\frac{r}{k_{i}}L$ and $\frac{r^{2}}{k_{i}^{2}}L$, which directly influence the convergence behavior. Further details are presented in Appendix F.3.
Based on the results in Theorem 4.4, we obtain the following corollary by applying an appropriate learning rate $\gamma$ to Algorithm 1. The proof can be found in Appendix F.4.
Corollary 4.5.
Under the same assumptions of Theorem 4.4, let $\mathcal{F}_{0}=f^{\mathcal{S}}({\bf X}^{0})-f^{*}$. Then there exists a learning rate $\gamma$ meeting the condition outlined in Theorem 4.4 such that the iterates $\{{\bf X}^{t}\}_{t=0}^{T-1}$ generated by FSLoRA satisfy
$\frac{1}{T}\sum_{t=0}^{T-1}\!\mathbb{E}\left\|\nabla_{\bf X}f^{\mathcal{S}}({\bf X}^{t})\right\|^{2}\!\leq\!8\frac{\sqrt{\bar{L}\mathcal{F}_{0}\sigma_{\rho}^{2}}}{\sqrt{NTH}}+10(\widetilde{L}L)^{\frac{1}{3}}\left(\frac{\mathcal{F}_{0}\sigma_{\rho}}{T}\right)^{\frac{2}{3}}+\frac{4\mathcal{F}_{0}}{T}.$ (7)
Discussion: Corollary 4.5 establishes an upper bound on the convergence of the proposed FSLoRA algorithm. The parameters $\bar{L}$ and $\widetilde{L}$ provide insight into how the sketching operation influences the convergence rate. Increasing $k_{i}$ would lead to a faster convergence for FSLoRA. However, this comes at the cost of increased communication and computational overhead for device $i$, indicating a trade-off in the selection of the sketching ratios. Additionally, the upper bound vanishes as $T\rightarrow\infty$. Moreover, the rate at which the bound diminishes is dominated by the first term, which recovers the convergence behavior of FedAvg [41, 19, 18] as the sketching ratio $k_{i}/r\to 1$(i.e., $\bar{L}=L$). This highlights the tightness of our analysis and shows that FSLoRA retains the convergence guarantees of vanilla Federated LoRA in the limit.
@@ 5 Experiments
Our experiments focus on RoBERTa (125M) [27] and LLaMA-3.2-3B [9], which represent typical model sizes suitable for on-device deployment, as well as the LLaMA-7B model to reflect large-scale scenarios. For RoBERTa and LLaMA-3.2-3B models, we fine-tune and evaluate them on the GLUE [35] and commonsense reasoning benchmark [16], respectively. For the LLaMA-7B model, we utilize Wizard, Dolly-15k, and Alpaca datasets, where the results are reported in Appendix D. Similar to [43, 37], we adopt Dirichlet-based partitioning for dataset splits. All the experiments are conducted on a cluster equipped with 4 NVIDIA A100 GPUs, each with 40 GB of memory. The number of devices is set to $20$ in the main manuscript, and to $50$ and $100$ in Appendix C. Further details are provided in the Appendix G.
@@ 5.1 Main Results Under Heterogeneous LoRA Setup
Performance comparison with baselines: We consider three state-of-the-art baselines listed in Section 2.2. For FSLoRA, the rank of the global LoRA modules is fixed as $r=64$, while the sketching ratio for device $i$ is sampled from the set $\{0.125,0.25,0.5,0.75\}$. For a fair comparison, we apply the same rank configuration to all baseline methods. Table 5.1 presents the performance of FSLoRA and baseline methods. Across both settings, FSLoRA consistently achieves superior accuracy while maintaining low GPU hours. In the GLUE & RoBERTa task, FSLoRA outperforms all baselines on average, with significant gains in MRPC, CoLA, and MNLI. In the commonsense reasoning & LLaMA task, which introduces higher model complexity, FSLoRA also delivers the best overall performance. Notably, FSLoRA achieves this while preserving computational efficiency comparable to HeteroLoRA as reflected in GPU hours. These results highlight FSLoRA’s effectiveness and scalability in heterogeneous LoRA fine-tuning scenarios.
Evaluation under broader heterogeneity, increased number of devices, and larger model: In Appendix C, we extend our evaluation to $50$ and $100$ devices, incorporating greater diversity in devices’ communication and computation capabilities, as well as varying levels of data heterogeneity. In Appendix D, we further assess the effectiveness of our method on the LLaMA-7B model.

### v3

@@ 3.2 Sparsity in the Gradients
In this subsection, we analyze the gradient structure of LoRA modules and highlight the gradients’ sparsity properties under a given sketching matrix. To begin, we present the gradient expressions for the LoRA modules $\bf B$ and $\bf A$ in the following lemma. The proof is provided in Appendix J.2.
@@
In particular, a random-$k$ diagonal sketching matrix selectively samples $k$ rows or columns of a matrix through left product or right product, respectively. With $\bf S$} being a random-$k$ diagonal matrix, the gradients of $\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$ with respect to LoRA modules $\bf B$ and $\bf A$, as shown in (3), naturally become structurally sparse matrices. This sparsity reduces computational and memory overhead during training, enabling faster gradient computation and parameter updates, while alleviating communication overhead across distributed clients by transmitting only non-zero elements.
@@
A key advantage of our formulation is its flexible control over the sparsity level of local gradients, achieved by configuring the parameter $k_{i}$ of the sketching matrix set $\mathcal{S}_{i}=\mathcal{S}(r,k_{i})$. This mechanism allows each client to tailor its local updates according to its communication and computation resource constraints, ensuring efficient and scalable fine-tuning in heterogeneous federated systems. Lowering $k_{i}$ helps resource-constrained clients reduce computation and communication overhead, while more capable clients can increase $k_{i}$ to conduct more informative local updates. Additionally, the distinction in sparsity level control between the proposed FSLoRA and the FedBCGD algorithm [28] is elaborated in Appendix B.
Remark 3.4 (Justification for the Choice of Random-$k$ Sketching).
We adopt Random-$k$ sketching due to its unbiasedness and the structured sparsity it induces. A detailed discussion and empirical comparison with alternative sketching strategies are provided in Appendix C.
@@ 3.3 FSLoRA Algorithm
Based on the formulation in (2), we propose a resource-adaptive algorithm termed FSLoRA for collaborative LLM fine-tuning. FSLoRA allows each client to update submatrices of the original modules $\bf{B}$ and $\bf{A}$ in each round. The server maintains a pair of global LoRA modules $\mathbf{B}$ and $\mathbf{A}$ and periodically updates them by aggregating sparse local updates received from distributed clients. Specifically, the procedure of FSLoRA at each round is detailed below.
  • The server begins by sampling sketching matrices $\{{\bf S}_{i}^{t}\sim\mathcal{S}_{i}\}_{i=1}^{N}$ for all clients, where $\mathcal{S}_{i}$ represents the set of possible sketching matrices for client $i$. These sketches are then sent to the corresponding clients. Additionally, the server broadcasts the current global LoRA modules $\mathbf{B}^{t};\mathbf{A}^{t}]$ to all clients. Note that the communication load introduced by transmitting the sketching matrix is negligible compared to global LoRA modules, as it involves only binary sketching indices (i.e., the diagonal elements of the sketching matrix); see Appendix A for details.
  • Clients perform local fine-tuning using sketch ${\bf S}_{i}^{t}$. Specifically, guided by sketching matrix ${\bf S}_{i}^{t}$, the update at client $i$ during the $h$-th iteration of the $t$-th round is given by:
@@
    where $\gamma$ denotes the learning rate and ${\Delta}{\bf B}_{i}^{t,h};{\Delta}{\bf A}_{i}^{t,h}]$ is a shorthand representation for:
@@
    The update direction in (4) corresponds to the negative stochastic gradient of $\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$ with respect to ${\bf B};{\bf A}]$ for a given sketch ${\bf S}_{i}^{t}$, as established in Lemma 3.2. The total update for client $i$ during one round of training, consisting of $H$ local steps, can be expressed as follows:
@@
    From the above equation, we see that only the columns of $\mathbf{B}$ and the rows of $\mathbf{A}$ corresponding to the nonzero entries of $\mathbf{S}_{i}^{t}$ are updated during the $t$-th round at client $i$. In essence, $\mathbf{S}_{i}^{t}$ selectively activates specific columns of $\mathbf{B}$ and rows of $\mathbf{A}$ for each round. Afterward, clients transmit these nonzero columns and rows of the sparse model updates to the server.
@@
The overall procedure of FSLoRA is summarized in Algorithm 1.
@@
Existing works on federated LoRA primarily adopt two aggregation strategies: (1) aggregating the LoRA modules as $\mathbf{B};\mathbf{A}]$ (e.g., vanilla Federated LoRA [47]), and (2) aggregating the product $\mathbf{B}\mathbf{A}$ (e.g., FlexLoRA [1]). Both methods have demonstrated effectiveness, as evidenced by their promising performance in prior studies. In this work, we adopt the former, as it introduces minimal overhead and retains the simplicity of LoRA. Additionally, we establish the convergence of FSLoRA under this aggregation choice in Section 4. We also demonstrate that FSLoRA is compatible with secure aggregation in Appendix D.
@@
The proposed FSLoRA introduces no additional operations compared to the vanilla Federated LoRA [47], resulting in minimal overhead for both the server and the clients relative to other heterogeneous LoRA baselines [1, 42]. A more detailed comparison of computation, memory, and communication is provided in Appendix A.
@@
2:  Server samples sketching matrices $\{{\bf S}_{i}^{t}\sim\mathcal{S}_{i}\}_{i=1}^{N}$ and sends them back to the clients
3:  Server broadcasts the current global LoRA modules to the clients
@@
5:   Clients update the local LoRA modules via (4)
@@
7:  Clients upload the non-zero columns of $({\bf B}_{i}^{t,H}\!-\!{\bf B}_{i}^{t,0})$ and the non-zero rows $({\bf A}_{i}^{t,H}\!-\!{\bf A}_{i}^{t,0})$
@@ 3.4 Comparison with Communication Compression
Although both the sketching approach in FSLoRA and communication compression [25] reduce communication overhead, the sketching approach fundamentally differs from traditional compression techniques. Compression methods focus solely on reducing the transmission load, leaving the gradient computation and model updates unchanged from the vanilla Federated LoRA. FSLoRA goes beyond communication savings by also reducing gradient computation and model update overhead through sparse training. Notably, these two methods are orthogonal and can be combined to achieve greater efficiency. Specifically, the compression can be applied to the transmission of non-zero columns of $\mathbf{B}$ and the non-zero rows of $\mathbf{A}$ in FSLoRA to further enhance communication efficiency. We demonstrate the effectiveness of this combination in Appendix H.4.
@@ 4 Analysis
In this section, we analyze the convergence of the proposed FSLoRA algorithm. We show that the iterate sequence generated by the FSLoRA algorithm converges to a stationary point of the function (2). Our analysis relies on the following notations.
Notations: We define $\tilde{\ell}({\bf B},{\bf A},\xi;{\bf S})=\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$ and $\tilde{f}_{i}({\bf B},{\bf A};{\bf S})=\mathbb{E}_{\xi\sim\mathcal{D}_{i}}\left[\ell({\bf W}_{0}+{\bf B}\bf S{\bf A},\xi)\right]$ for a given $\bf S$} and $f_{i}^{\mathcal{S}}(\mathbf{B},\mathbf{A})=\mathbb{E}_{{\bf S}\sim\mathcal{S}_{i}}[\tilde{f}_{i}({\bf B},\mathbf{A};{\bf S})]$. For simplicity, we denote $\mathbf{X}=[\mathbf{B};\mathbf{A}]$ and rewrite $f(\mathbf{B},\mathbf{A})$, $f_{i}(\mathbf{B},\mathbf{A})$, $f^{\mathcal{S}}(\mathbf{B},\mathbf{A})$, $f_{i}^{\mathcal{S}}(\mathbf{B},\mathbf{A})$, $\tilde{f}_{i}({\bf B},{\bf A};{\bf S})$, and $\tilde{\ell}({\bf B},{\bf A},\xi;{\bf S})$ as $f(\mathbf{X})$, $f_{i}(\mathbf{X})$, $f^{\mathcal{S}}(\mathbf{X})$, $f_{i}^{\mathcal{S}}(\mathbf{X})$, $\tilde{f}_{i}(\mathbf{X};\mathbf{S})$, and $\tilde{\ell}({\bf X},\xi;{\bf S})$ respectively. In addition, we use $\|\cdot\|$ to denote the Frobenius norm.
@@
Assumption 4.1 is standard in optimization literature [3, 15, 4]. Assumptions 4.2 and 4.3 are commonly adopted in federated learning to bound sampling randomness and data heterogeneity [14, 44]. We further provide an empirical validation in Appendix E, showing that Assumptions 4.2 and 4.3 are reasonable within the LLM fine-tuning scenario. Building on these assumptions, we analyze the convergence behavior of FSLoRA. Our main results are summarized in the following theorem.
@@
Suppose that Assumptions 4.1-4.3 hold, then there exists a learning rate $\gamma\leq\min\{\frac{N}{24\rho(c_{h}+1)H\bar{L}},\frac{1}{8\sqrt{\widetilde{L}L(\rho+1)(c_{h}+1)}H},\frac{1}{H}\}$ such that the iterates $\{{\bf X}^{t}\}$ generated by FSLoRA satisfy
$\frac{1}{T}\sum_{t=0}^{T-1}\!\mathbb{E}\left\|\nabla_{\bf X}f^{\mathcal{S}}({\bf X}^{t})\right\|^{2}\!\leq\!8\frac{\sqrt{\bar{L}\mathcal{F}_{0}\sigma_{\rho}^{2}}}{\sqrt{NTH}}+10(\widetilde{L}L)^{\frac{1}{3}}\left(\frac{\mathcal{F}_{0}\sigma_{\rho}}{T}\right)^{\frac{2}{3}}+\frac{4\mathcal{F}_{0}}{T},$ (6)
where $\sigma_{\rho}^{2}=\sigma^{2}+3(\rho+1)\sigma_{h}^{2}$, $\bar{L}=\left(\frac{1}{N}\sum_{i=1}^{N}\frac{r}{k_{i}}\right)L$, $\widetilde{L}=\left(\frac{1}{N}\sum_{i=1}^{N}\frac{r^{2}}{k^{2}_{i}}\right)L$ and $\mathcal{F}_{0}=f^{\mathcal{S}}({\bf X}^{0})-f^{*}$ with $f^{*}$ denoting the lower bound of $f^{\mathcal{S}}({\bf X})$.
Technical highlights of Theorem 4.4: A key step in the proof of Theorem 4.4 is characterizing the impact of the sketching mechanism on the optimization landscape. Our analysis reveals how the sketching operation modifies the smoothness properties of the objective, introducing scaled smoothness constants, $\frac{r}{k_{i}}L$ and $\frac{r^{2}}{k_{i}^{2}}L$, which directly influence the convergence behavior. Further details are presented in Appendix J.3.
Discussion: Theorem 4.4 establishes an upper bound on the convergence of the proposed FSLoRA algorithm. The parameters $\bar{L}$ and $\widetilde{L}$ provide insight into how the sketching operation influences the convergence rate. Increasing $k_{i}$ would lead to a faster convergence for FSLoRA. However, this comes at the cost of increased communication and computational overhead for client $i$, indicating a trade-off in the selection of the sketching ratios. Additionally, the upper bound vanishes as $T\rightarrow\infty$. Moreover, the rate at which the bound diminishes is dominated by the first term, which recovers the convergence behavior of FedAvg [45, 22, 21] as the sketching ratio $k_{i}/r\to 1$(i.e., $\bar{L}=L$). This highlights the tightness of our analysis and shows that FSLoRA retains the convergence guarantees of vanilla Federated LoRA in the limit.
@@ 5 Experiments
Our experiments focus on RoBERTa (125M) [29] and LLaMA-3.2-3B [12], which represent typical model sizes suitable for client-side deployment, as well as the LLaMA-7B model to reflect large-scale scenarios. For RoBERTa and LLaMA-3.2-3B models, we fine-tune and evaluate them on the GLUE [39] and commonsense reasoning benchmark [19], respectively. For the LLaMA-7B model, we utilize Wizard, Dolly-15k, and Alpaca datasets, where the results are reported in Appendix G. Similar to [47, 42], we adopt Dirichlet-based partitioning for dataset splits. All the experiments are conducted on a cluster equipped with 4 NVIDIA A100 GPUs, each with 40 GB of memory. The number of clients is set to $20$ in the main manuscript, and to $50$ and $100$ in Appendix F. Further details are provided in the Appendix I. The implementation code for this project is included in the supplementary material.
@@ 5.1 Main Results Under Heterogeneous LoRA Setup
Performance comparison with baselines: We consider three state-of-the-art baselines listed in Section 2.2. For FSLoRA, the rank of the global LoRA modules is fixed as $r=64$, while the sketching ratio for client $i$ is sampled from the set $\{0.125,0.25,0.5,0.75\}$. For a fair comparison, we apply the same rank configuration to all baseline methods. Table 5.1 presents the performance of FSLoRA and baseline methods. Across both settings, FSLoRA consistently achieves superior accuracy while maintaining low GPU hours. In the GLUE & RoBERTa task, FSLoRA outperforms all baselines on average, with significant gains in MRPC, CoLA, and MNLI. In the commonsense reasoning & LLaMA task, which introduces higher model complexity, FSLoRA also delivers the best overall performance. Notably, FSLoRA achieves this while preserving computational efficiency comparable to HeteroLoRA as reflected in GPU hours. These results highlight FSLoRA’s effectiveness and scalability in heterogeneous LoRA fine-tuning scenarios.
Evaluation under broader heterogeneity, increased number of clients, and larger model: In Appendix F, we extend our evaluation to $50$ and $100$ clients, incorporating greater diversity in clients’ communication and computation capabilities, as well as varying levels of data heterogeneity. In Appendix G, we further assess the effectiveness of our method on the LLaMA-7B model.